# Nightjar: Dynamic Adaptive Speculative Decoding for Large Language Models Serving

### v2

@@ 6.3.2 Efficacy of Dynamic Model Offloading
The impact of our dynamic loading and unloading mechanism is assessed by comparing Nightjar-Offload with the standard Nightjar implementation. Results in Figure 13 demonstrate that offloading becomes increasingly beneficial as the system enters the compute-bound regime at high request rates, reaching a peak throughput of 6315.9 tok/s at 35 req/s—a substantial improvement over the 5982.6 tok/s achieved without offloading. This confirms that liberating GPU memory by unloading draft models effectively expands the KV cache capacity for the target model, directly translating to higher service concurrency.
@@ 7.1 Speculative Decoding in LLM Serving
While various approaches have sought to optimize speculative decoding, they often fail to address the dynamic and unpredictable nature of real-world LLM serving loads. Initial efforts focused on single-request optimization [26, 31, 32, 6, 11] or static offline batches [25, 10], employing predictive algorithms [31] or confidence-based methods [12, 26, 27], but their static nature makes them fundamentally ill-suited for the fluctuating request rates in live environments. More recent works have explored dynamic speculation length adjustment in serving systems [25, 19, 12, 18], yet they exhibit significant limitations. For instance, SpecServe [12] and TETRIS [27] rely on draft confidence scores, but this can lead to poor performance with large batch sizes and still requires generating at least one token. DSD [19] optimizes length via linear goodput modeling, but its reliance on past average acceptance rates introduces a critical vulnerability. This dependency on historical data, found in methods like DSD [19] and SpecServe [12], can lead to a “decision deadlock” where disabling speculative decoding permanently halts the collection of new acceptance rate data, locking the system in a suboptimal state. Our exploration-exploitation approach is designed to mitigate this risk. Even the work of BanditSpec [10], which first introduced a Multi-Armed Bandit (MAB) paradigm, is constrained by its static design, rendering it incompatible with the continuous batching [13] of modern serving systems. Furthermore, to the best of our knowledge, existing works largely overlook the “switching cost” incurred when re-enabling speculative decoding from a length of 0, which is the overhead of KV cache reconstruction for the draft model.
@@ 7.2 Unified Memory Management for Weights and KV Cache in LLM Serving
Oneiros [17] targets multi-model deployment by introducing a dynamic memory remapping engine that redirects GPU memory reserved for model parameters to the KV cache. Unlike Nightjar, which completely unloads draft model weights when they are not in use , Oneiros operates on active models. By performing remapping at layer granularity, it enables models to remain in a serving state while utilizing asynchronous transfers to fetch required parameters in real-time before execution. This approach effectively hides latency by overlapping computation with parameter transfer. eLLM [29] focuses on intra-model efficiency, enabling the KV cache and activation tensors to ”borrow” space from a shared physical memory pool, effectively breaking the traditional isolation between these resources. FineServe [5] provides a unified management framework for both weights and KV caches across multiple models, specifically addressing the challenges of serving models with heterogeneous quantization precisions.

### v3

@@ 7.3.2 Efficacy of Dynamic Model Offloading
For 7B model, the impact of our dynamic loading and unloading mechanism is assessed by comparing Nightjar-Offload with the standard Nightjar implementation. Results in Figure 13 demonstrate that offloading becomes increasingly beneficial as the system enters the compute-bound regime at high request rates, reaching a peak throughput of 6315.9 tok/s at 35 req/s, a substantial improvement over the 5982.6 tok/s achieved without offloading. This confirms that liberating GPU memory by unloading draft models effectively expands the KV cache capacity for the target model, directly translating to higher service concurrency.
@@ 8.1 Speculative Decoding in LLM Serving
While various approaches have sought to optimize speculative decoding, they often fail to address the dynamic and unpredictable nature of real-world LLM serving loads. Initial efforts focused on single-request optimization [28, 33, 34, 6, 12] or static offline batches [27, 11], employing predictive algorithms [33] or confidence-based methods [13, 28, 29], but their static nature makes them fundamentally ill-suited for the fluctuating request rates in live environments. More recent works have explored dynamic speculation length adjustment in serving systems [27, 21, 13, 20], yet they exhibit significant limitations. For instance, SpecServe [13] and TETRIS [29] rely on draft confidence scores, but this can lead to poor performance with large batch sizes and still requires generating at least one token. DSD [21] optimizes length via linear goodput modeling, but its reliance on past average acceptance rates introduces a critical vulnerability. This dependency on historical data, found in methods like DSD [21] and SpecServe [13], can lead to a “decision deadlock” where disabling speculative decoding permanently halts the collection of new acceptance rate data, locking the system in a suboptimal state. Our exploration-exploitation approach is designed to mitigate this risk. Although BanditSpec [11] introduces a Multi-Armed Bandit (MAB) paradigm, it is constrained by its static design, rendering it incompatible with the continuous batching [14] of modern serving systems. Furthermore, to the best of our knowledge, existing works largely overlook the “switching cost” incurred when re-enabling speculative decoding from a length of 0, which is the overhead of KV cache reconstruction for the draft model.
@@ 8.2 Unified Memory Management for Weights and KV Cache in LLM Serving
Oneiros [18] targets multi-model deployment by introducing a dynamic memory remapping engine that redirects GPU memory reserved for model parameters to the KV cache. Unlike Nightjar, which completely unloads draft model weights when they are not in use, Oneiros operates on active models. By performing remapping at layer granularity, it enables models to remain in a serving state while utilizing asynchronous transfers to fetch required parameters in real-time before execution. This approach effectively hides latency by overlapping computation with parameter transfer. eLLM [31] focuses on intra-model efficiency, enabling the KV cache and activation tensors to borrow space from a shared physical memory pool, effectively breaking the traditional isolation between these resources. FineServe [5] provides a unified management framework for both weights and KV caches across multiple models, specifically addressing the challenges of serving models with heterogeneous quantization precisions. While these works optimize memory sharing for active models or intra-model tensors, they do not address the unique memory bottlenecks associated with dynamically swapping draft model weights in speculative decoding, a gap our work aims to fill.
@@ 9 Conclusions
Appendix A Detailed Regret Analysis
In this appendix, we provide the complete derivation of the regret bound for the Nightjar algorithm. Our objective is to minimize the cumulative latency per token. A key aspect of our formulation is the specific structure of the loss function, which penalizes strategy switches inversely proportional to the speculative length.
A.1 Problem Setup and Definitions
Let $\mathcal{A}=\{0,1,\dots,\Gamma_{\max}\}$ be the set of available arms (speculative lengths). At each time step $t$, the algorithm selects an arm $\gamma_{t}\in\mathcal{A}$. We define the instantaneous Loss observed at step $t$ as:
$L_{t}(\gamma_{t})=\ell_{B,\gamma_{t}}+\mathbb{I}(\gamma_{t-1}=0\land\gamma_{t}>0)\cdot\frac{C_{\text{switch}}}{\gamma_{t}}$ (10)
where $\ell_{B,\gamma_{t}}=1/g_{B,\gamma_{t}}$ is the latency per token, and the second term represents the switching overhead amortized by the speculative length $\gamma_{t}$.
The Cumulative Regret over horizon $T$ is defined against an optimal fixed policy $\gamma^{*}$ that knows the true distributions:
$R(T)=\sum_{t=1}^{T}\left(L_{t}(\gamma_{t})-L_{t}(\gamma^{*})\right)$ (11)
Since the optimal policy $\gamma^{*}$ is static, it incurs the switching cost at most once (at $t=1$), which is negligible as $T\to\infty$. Thus, we can decompose the regret into performance regret and switching regret:
$R(T)\leq\underbrace{\sum_{t=1}^{T}(\ell_{B,\gamma_{t}}-\ell_{B,\gamma^{*}})}_{R_{\text{perf}}(T)}+\underbrace{\sum_{t=1}^{T}\mathbb{I}(\text{switch}_{t})\cdot\frac{C_{\text{switch}}}{\gamma_{t}}}_{R_{\text{switch}}(T)}$ (12)
A.2 Bound on Switching Cost Regret
We first analyze $R_{\text{switch}}(T)$. The specific design of our loss function scales the switching penalty by $1/\gamma_{t}$. Since any valid switch implies $\gamma_{t}\geq 1$, we have the inequality:
$\frac{C_{\text{switch}}}{\gamma_{t}}\leq C_{\text{switch}}$ (13)
This implies that the switching cost is bounded by a constant maximum overhead $C_{\max}$ regardless of the chosen arm.
The Nightjar algorithm employs a Bin Locking Mechanism, where the arm $\gamma_{t}$ is fixed for the duration of a bin. Therefore, a switch can only occur at the boundary of a bin. Let $S_{T}$ be the total number of bins initiated up to time $T$. As shown in the algorithm design, the timeline is divided into blocks $j$ of length $H_{j}=2^{j-1}$, and each block contains $K_{j}\approx\sqrt{H_{j}}$ bins.
Lemma 2 (Switching Count Bound).
The total number of strategy switches is bounded by the number of bins:
$N_{\text{bins}}\leq\sum_{j=1}^{\log_{2}T}\sqrt{2^{j-1}}=\mathcal{O}(\sqrt{T})$ (14)
Combining the cost bound and the count bound:
$R_{\text{switch}}(T)\leq C_{\max}\cdot N_{\text{bins}}=\mathcal{O}(C_{\max}\sqrt{T})=\tilde{\mathcal{O}}(\sqrt{T})$ (15)
This confirms that even with the inverse-proportional penalty term, the switching regret remains sublinear.
A.3 Bound on Performance Regret
The analysis for $R_{\text{perf}}(T)$ follows the standard decomposition into exploration and exploitation phases.
A.3.1 Exploration Regret
In block $j$, the algorithm explores with probability $p_{j,b}\propto 1/b$, where $b$ is the bin index. The expected number of exploration bins grows logarithmically with the block size:
$\mathbb{E}[N_{\text{explore}}^{(j)}]\approx\ln(\sqrt{H_{j}})$ (16)
Since each bin has duration $W_{j}\approx\sqrt{H_{j}}$, the regret from random exploration is:
$R_{\text{explore}}(T)=\sum_{j=1}^{\log_{2}T}\Delta_{\max}\sqrt{H_{j}}\ln\sqrt{H_{j}}=\tilde{\mathcal{O}}(\sqrt{T})$ (17)
A.3.2 Exploitation Regret
In exploitation bins, the algorithm selects the arm that minimizes the empirical estimate of the regularized loss defined in Eq. (10). Using the Hoeffding-Azuma inequality for martingales, the probability of selecting a suboptimal arm decays exponentially with the number of samples collected in previous bins. As established in [24], the exploitation regret is dominated by the exploration and switching terms, bounded by $\mathcal{O}(\sqrt{T})$.
A.4 Conclusion
Summing the components:
$R(T)=\tilde{\mathcal{O}}(\sqrt{T})+\tilde{\mathcal{O}}(\sqrt{T})=\tilde{\mathcal{O}}(\sqrt{T})$ (18)
The Nightjar algorithm achieves sublinear regret, effectively balancing the trade-off between maximizing goodput (minimizing latency) and minimizing the KV cache reconstruction overhead defined by the specific loss function $C_{\text{switch}}/\gamma$.